# Hidden long-range correlations in the ion distribution at the graphite / [bmim][NTf2] electrified interface

### v2

@@ IV Summary and discussion
The interface between ionic liquids and solid electrodes is a fascinating subject that joins the current interest for a vast number of electrochemical applications with concepts and theories dating back to one century ago.korn1 ; korn2 ; korn3 ; bock . This fertile combination, together with the complex structure of ions and the nanostructures texture of ionic liquids, has fostered a variety of results of practical interest, and also some surprises that challenge the accepted picture of the electrified interfaces. Among the aspects that are being challenged is the long held notion that, beyond a narrow interfacial region of microscopic width, a Coulombic fluid in contact with a charged interface is virtually bulk-like. A strong indication in this direction is provided by the recent observationexpLangmuir of the Pockel effect at the interface between the [bmim][NTf2] ionic liquid and two planar solid parallel electrodes connected to a generator maintaining an electrostatic potential energy difference between the two sides of the IL slab. Upon charging the interface, a change in the Raman intensity was observed, arising from all Raman-active modes of cations and anions, independently from their distance from the interface.
@@
To investigate this phenomenon, the interface between the same [bmim][NTf2] ionic liquid and a graphite electrode has been simulated by molecular dynamics, using an atomistic semi-empirical force field. An innovative analysis of trajectories allows to join [bmim]+ and [NTf2]- ions into neutral pairs, whose dipoles show an apparent alignment with the external electric field driven by the charge deposited on the graphite surface. This alignment is observed even well inside the IL bulk-like region, where the external field is exactly cancelled (on average) by the screening at the interface by the mobile ions in the IL. Since dipoles do not obey the strict screening conditions of charges, the observation does not violate any accepted theory of electrified interfaces, but opens the way to a whole new class of medium-long range correlations that remain hidden when one considers the system only as a collection of free ions.
@@
Up to now, this observation concerns a geometric property in the distribution of ions throughout the system, that leaves open questions concerning its origin and its significance with respect to the observed effect on the Raman spectra measured on the experimental samples. The main question on its origin concerns the role of the shape anisotropy of the ions in transferring the information on the state of the interface deep into the IL bulk. The main question on the relevance of the geometric property concerns whether this correlation among virtual dipoles can alter the vibrational properties of the ionic liquid. The mechanism joining the geometric correlation with the change of Raman spectroscopic properties has to respect the fact that the average electric field well inside the ionic liquid vanishes, independently from the charge on the electrode. This fact, however, does not prevent changes in the average square amplitude of the fluctuating electric field inside the ionic liquid, that could indeed give origin to the observed change of vibrational frequencies.
@@
Because of these open questions that are currently being investigated, the present document represents a report on work in progress, that will be followed by a full manuscriprit on the comprehensive investigation of the [bmim][NTf2] / graphite interface and on the hidden correlations in the distribution of ions arising upon charging the electrodes.
@@
Further considerations that will be explored during later stages of our investigations are as follows. Long range correlations imply that the two electrodes in simulated planar capacitors are likely to interact up to the largest sizes that are practically achievable right now, affecting any estimate of properties that depend on the spatial distribution and orientation of the ions. This problem enhances the appeal of single-electrode schemes like the one introduced in Ref. eletransf, , in which a unique spherical electrode surrounds the pool of electrolyte, whose thermodynamic state is controlled by adopting the grand canonical simulation approach.

### v3

@@ IV Summary and discussion
The interface between ionic liquids and solid electrodes is a fascinating subject that joins the current interest for a vast number of electrochemical applications with concepts and theories dating back to one century ago.korn1 ; korn2 ; korn3 ; bock . This fertile combination, together with the complex structure of ions and the nanostructured texture of ionic liquids, has fostered a variety of results of practical interest, and also offered genuine surprises that challenge the accepted picture of the electrified interfaces. Among the aspects that are being challenged is the long held notion that, beyond a narrow interfacial region of microscopic width, a Coulombic fluid in contact with a charged interface is virtually bulk-like. A strong indication in this direction is provided by the recent observationexpLangmuir of the Pockel effect at the interface between the [bmim][NTf2] ionic liquid and two planar solid parallel electrodes connected to a generator maintaining an electrostatic potential energy difference between the two sides of the IL slab. Upon charging the interface, a change in the Raman intensity was observed, arising from all Raman-active modes of cations and anions, independently from their distance from the interface.
@@
To investigate this phenomenon, the interface between the same [bmim][NTf2] ionic liquid and a graphite electrode has been simulated by molecular dynamics, using an atomistic semi-empirical force field. An innovative analysis of trajectories allows to join [bmim]+ and [NTf2]- ions into neutral pairs, whose dipoles show an apparent alignment with the external electric field driven by the charge deposited on the graphite surface. This alignment is observed even well inside the IL bulk-like region, where the external field is exactly cancelled (on average) by the screening at the interface by the mobile ions in the IL. Since dipoles do not obey the same strict screening conditions of charges, the observation does not violate any accepted theory of electrified interfaces, but opens the way to a whole new class of medium-long range correlations that remain hidden when one considers the system only as a collection of free ions.
@@
Up to now, this observation concerns a geometric property in the distribution of ions throughout the system, that leaves open questions concerning its origin and its significance with respect to the observed effect on the Raman spectra measured on the experimental samples. The main question on its origin concerns the role of the shape anisotropy of the ions in transferring the information on the state of the interface deep into the IL bulk. The remaining question on the relevance of the geometric property concerns whether this correlation among virtual dipoles can alter the vibrational properties of the ionic liquid. The mechanism joining the geometric correlation with the change of Raman spectroscopic properties has to respect the fact that the average electric field well inside the ionic liquid vanishes, independently from the charge on the electrode. This fact, however, does not prevent changes in the average square amplitude of the fluctuating electric field inside the ionic liquid, that could indeed give origin to the observed change of vibrational frequencies. A preliminary investigation of this scenario by MD did not provide a definite answer, since the effect is likely to be small and acquiring sufficient statistics requires a very intensive effort.
@@
Because of these open questions that are currently being investigated, the present document represents a report on work in progress, that will be followed by a full manuscript on the comprehensive investigation of the [bmim][NTf2] / graphite interface and on the hidden correlations in the distribution of ions arising upon charging the electrodes.
@@
Further considerations that will be explored during later stages of our investigations are as follows. Long range correlations imply that the two electrodes in simulated planar capacitors are likely to interact up to the largest sizes that are practically achievable right now by computer simulation, affecting any estimate of properties that depend on the spatial distribution and orientation of the ions. This problem enhances the appeal of single-electrode schemes like the one introduced in Ref. eletransf, , in which a unique spherical electrode surrounds the pool of electrolyte, whose thermodynamic state is controlled by adopting the grand canonical simulation approach.
@@
As a last remark, we observe that our geometric construction building (instantaneous) dipoles out of a population of mobile ions has some points of contact with the derivation in Ref. bazant, of a spin-glass Hamiltonian mimicking structural properties of viscous ionic liquids, which is then used to characterize a variety of correlation properties. The difference is that, while Ref. bazant, emphasizes many-body concepts, the present approach closely follows traditional lines of electrostatics, defining new entities (dipoles) from simpler components like ions. This last procedure makes sense in this case, in which ions are complex, their motion is sluggish at best, and combining them into dipoles decreases the range of their interaction, opening the way to the types of correlations discussed in the present paper.